OpenStudy (anonymous):

Find an equation for the tangent plane to the surface given by Υ(u; v) = (u^2- v^2; u + v; u^2 + 3v) at the point (3; 1; 1).

OpenStudy (amistre64):

looks like you might need to use the chain rule ...

OpenStudy (amistre64):

F = (x,y,z) x = u^2 + v^2 x' = 2u u' + 2v v' y = u+v y' = u' + v' z = u^2 + 3v z' = 2u u' + 3v'

OpenStudy (amistre64):

solve the system ... u^2- v^2 = 3 u + v = 1 u^2 + 3v = 1 maybe ....

OpenStudy (amistre64):

\[\frac{df}{du}=\frac{df}{dx}\frac{dx}{du}+\frac{df}{dy}\frac{dy}{du}+\frac{df}{dz}\frac{dz}{du}\] \[\frac{df}{du}=\nabla f\cdot <x'(u),y'(u)>,z'(u)\]

OpenStudy (amistre64):

...latex fail

OpenStudy (amistre64):

regardless x = u^2 + v^2 x'(u) = 2u x'(v) = 2v y = u+v y'(u) = 1 y'(v) = 1 z = u^2 + 3v z'(u) = 2u z'(v) = 3 gradF = <1,1,1> Fu = 4u+1 Fv = 2v+4